OpenStudy (anonymous):

Is the answer to sqrt x + sqrt x-3=3 11/2?

OpenStudy (anonymous):

To check you renter your solution back into the original equation \[\sqrt{\frac{ 11 }{ 2 }}+\sqrt{\frac{ 11 }{ 2 }-3}=3.926...\] which is not 3

OpenStudy (anonymous):

sorry top line was suppose to say: To check your solution renter it back into the original equation

OpenStudy (anonymous):

Where did I go wrong? I screwed up the problem somehow.

OpenStudy (anonymous):

well use the equation editor to show me the steps you took to solve the problem

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{x-3}=3\]

OpenStudy (anonymous):

\[\left( \sqrt{x} \right)^{2}+(\sqrt{x-3})^{2}=\left( 3 \right)^{2}\]

OpenStudy (anonymous):

\[x+x-3=9\]

OpenStudy (anonymous):

stop that is where you went wrong

OpenStudy (anonymous):

\[2x-3=9\]

OpenStudy (anonymous):

Okay. What do I do?

OpenStudy (anonymous):

you are squaring the entire side \[(\sqrt{x}+\sqrt{x-3})^{2}=3^{2}\] not your earlier step but I think this will give you more problems then it will solve give me a second and I will see if I can find an easier path t lead you down

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{x-3}=3\] is this your equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It might help me if you tell me what the title of the section this question is under is.

OpenStudy (anonymous):

Equations with radicals. It just asks to solve the following equations.

OpenStudy (anonymous):

what techniques have you been taught?

OpenStudy (anonymous):

Well, when it comes to problems like this, all I've learned is to square both sides and go from there but apparently, I'm not very good at it.

OpenStudy (anonymous):

I need to close this one to find help with another one but if you could continue helping me with this one, that would be great! :)

OpenStudy (anonymous):

I know the answer is 4 by graphing but I cannot see the algebraic method for some reason

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@hba this may be too easy for you but fr some reason I cannot see the algebraic way to solve this question so that I can lead him through it. It has been too long since I have had to solve radicals without calculus.